A component is assembled in a four-step process. The probability of making a defect in step 1 is 10%; in step 2, 20%; in step 3, 8%; and in step 4, 5%. What is the probability that the component will be defective?

99.99%

62.9%

37.1%

0.008%
Answer #1

Correct answer is 37.1 % . Explaination is given below :

Production Probability % at each step for producing a defective piece Probability % at each step for producing a defect free piece Probability of producing a correct piece Probability of producing a defective piece
steps deffective without defects 100 100-100=0
1 10 90 100*90/100=90 100-90=10
2 20 80 90*80/100=72 100-72=28
3 8 92 72*92/100=66.24 100-66.24-33.76
4 5 95 66.24*95/100=62.928 100-62.928=37.1